11. How much aluminum can be produced from the \(1.00 \times 10^{3} \mathrm{~kg}\) mineral bauxite, which mostly \(\mathrm{Al}_{2} \mathrm{O}_{3} ?\)
Answer #1

Bauxite = Al2O3

The balanced reaction of production of Al from Al2O3

2Al2O3 = 4Al + 3O2

Mass of Al2O3 = 1.00*10^3 kg

Moles of Al2O3 = mass/molecular weight

= (1.00*10^3 kg) / (101.96 kg/kmol)

= 9.808 kmol

From the stoichiometry of the reaction

2 kmoles Al2O3 produces = 4 kmoles Al

9.808 kmol Al2O3 produces = 9.808*4/2 kmoles Al

= 19.616 kmol

Mass of Al = moles x molecular weight

= 19.616 kmol x 26.98 kg/kmol

= 529.23 kg

= 5.29 * 10^2 kg
Answer #2

Q11. Given : mass Al2O3 = 1.00 x 103 kg

mass Al2O3 = 1.00 x 103 kg * (1000 g / 1 kg)

mass Al2O3 = 1.00 x 106 g

moles Al2O3 = (mass Al2O3) / (molar mass Al2O3)

moles Al2O3 = (1.00 x 106 g) / (102 g/mol)

moles Al2O3 = 9803.9 mol

Al2O3 (s) \rightarrow 2 Al (s) + 3/2 O2 (g)

moles Al produced = 2 * (moles Al2O3)

moles Al produced = 2 * (9803.9 mol)

moles Al produced = 19607.8 mol

mass Al produced = (moles Al produced) * (molar mass Al)

mass Al produced = (19607.8 mol) * (27.0 g/mol)

mass Al produced = 529411.8 g

mass Al produced = 529411.8 g * (1 kg / 1000 g)

mass Al produced = 529 kg
Answer #4

The boonud lachen 15 2 A10,4 - HAL(9) + 302 Moloy mag of Alb - 10/96 g/mol solo aromic mass of AL -26.989/mol QY 101.969 of A
Answer #3

moles of Al₂O3= mass molar mass of Al2O3 3 = 1X109 xlog - 9.81 X10% mol 101.96 gimos 10 Now, Imole Al2O3 contains a 2 mol Al